OpenStudy (anonymous):

Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times as many units of plywood as lumber. It makes a profit of $30 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $18540?

OpenStudy (anonymous):

let x=number of units of lumber; p=number of units of plywood then p=3x 30x+50p=18540 replace p=3x to get 30x+50*3x=18540 solve for x 180x=18540 x=103 units of lumber and p=3*103=309 units of plywood Check: 30(103)+50(309)=18540